OpenStudy (anonymous):

Factor r^3 + 3r^2 -10r -24

OpenStudy (anonymous):

\[r^3 + 3r^2 - 10r - 24\] Factored out the common factor of the polynomial: \[(r^2 + 4 - 12)(r+2) \] Factored the quadratic into a binomial (answer): \[(r−3)(r+4)(r+2)\]

OpenStudy (anonymous):

Can there also be different forms of this, like other factors?

OpenStudy (anonymous):

And I am not sure what you mean by the common factor sorry...

OpenStudy (anonymous):

Okay, first of all I made a typo in the second equation the 4 is suppose to be an r

OpenStudy (anonymous):

\[r^3 + 3r^2 - 10r - 24\]

OpenStudy (anonymous):

I think you were missing a factor. (r-1)(r+2)(r+4)(r-3)

OpenStudy (anonymous):

The first thing to do is find the common factor, do you know what greatest common factor is?

OpenStudy (anonymous):

I am just looking for some insight of where to start on these. My professor told me that when the coefficients add up to zero I can use r-1 as a factor

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But i dont see how r-2 goes into all those terms. I have never really looked for gcf in this way before unfortunately.

OpenStudy (anonymous):

Okay, well in this equation first we want to find the factors of 24, and see which factor makes the equation equal to 0 when we substitute it for r.

OpenStudy (anonymous):

ahh ok

OpenStudy (anonymous):

So all the solutions are going to be factors of the last term?

OpenStudy (anonymous):

What we should be left with is a factor beside a quadratic which can be simplified to 2 binomials. The first 2 expressions in the parenthesis is the quadratic simplified, and the third is actually the greatest common factor of the original equation.

OpenStudy (anonymous):

The factors of 24 are 1, 2, 3, 4, 6, 8, 12, and 24

OpenStudy (anonymous):

We try to see if when we input the positive or negative value of a factor if the equation equals 0

OpenStudy (anonymous):

And we find that when r = -2 is when the equation is equal to 0

OpenStudy (anonymous):

And I found that by testing 1, then -1, then 2, then -2, etc up the positive and negatives of the factors

OpenStudy (anonymous):

Basically because -2 made the equation equal to 0, (r+2) is a factor

OpenStudy (anonymous):

Thanks this really helped a lot (;

OpenStudy (anonymous):

No problem lol

hero (hero):

Applying the rational root theorem, r = -2 is a solution to the given expression and r + 2 = 0 is a factor. Next split given expression in accordance with factor x + 2 and factor completely: r^3 + 3r^2 -10r -24 = r^3 + 2r^2 + r^2 + 2r - 12r - 24 = r^2(r + 2) + r(r + 2) - 12(r + 2) = (r + 2)(r^2 + r - 12) = (r + 2)(r + 4)(r - 3)